You find yourself on a planet (Radius = 2.33 x 107m where a 2.00 kg mass has a weight of 32.0 N. What is the escape velocity from this planet ? 


1
Expert's answer
2021-05-18T19:46:55-0400

Given:

g(on this planet)= 16 m/s2

R = 2.33 x 107

and we have G = 6.67 × 10-11 Newtons kg-2 m2 as universal constant,


As we know the value of g on earth is given by

"g={GM \over R^2}"

and value of escape velocity on earth is given by


"V_s =\sqrt{2GM\over R}"

from both above equation we get

"V_s =\sqrt{2gR}"

now in this putting the given values of g and R of other planet we get,

"V_s =\sqrt{2(16)(2.33\times10^7)}"

"V_s=2.73\times10^4" m/s ...............answer